### v1

# CyC3D: Fine-grained Controllable 3D Generation via Cycle Consistency Regularization

Abstract
Despite the remarkable progress of 3D generation, achieving controllability, i.e., ensuring consistency between generated 3D content and input conditions like edge and depth, remains a significant challenge. Existing methods often struggle to maintain accurate alignment, leading to noticeable discrepancies. To address this issue, we propose CyC3D, a new framework that enhances controllable 3D generation by explicitly encouraging cyclic consistency between the second-order 3D content, generated based on extracted signals from the first-order generation, and its original input controls. Specifically, we employ an efficient feed-forward backbone that can generate a 3D object from an input condition and a text prompt. Given an initial viewpoint and a control signal, a novel view is rendered from the generated 3D content, from which the extracted condition is used to regenerate the 3D content. This re-generated output is then rendered back to the initial viewpoint, followed by another round of control signal extraction, forming a cyclic process with two consistency constraints. View consistency ensures coherence between the two generated 3D objects, measured by semantic similarity to accommodate generative diversity. Condition consistency aligns the final extracted signal with the original input control, preserving structural or geometric details throughout the process. Extensive experiments on popular benchmarks demonstrate that CyC3D significantly improves controllability, especially for fine-grained details, outperforming existing methods across various conditions (e.g., +14.17% PSNR for edge, +6.26% PSNR for sketch).
1 Introduction
Figure 1: Visualization of input and reconstructed conditions. We feed the input condition to the generation model first and use the generated 3D contents to render the image at the input viewpoint to extract the reconstructed conditions. The difference between the input and the reconstructed conditions can reveal whether the fine-grained controls are reserved in the results.
Figure 2: Brief overview of our proposed CyC3D. We first feed the condition on the original reference view to the generation model and extract the conditions from the rendered results on randomly sampled novel views. The extracted conditions are then fed to the generation model again and rendered back to the original reference view. The extracted condition of the rendered results should be consistent with the input condition on the reference view.
The creation of 3D models from text descriptions (text-to-3D) or image collections (image-to-3D) is a fundamental task in computer graphics, attracting growing interest from a wide range of domains, including VR/AR [2], digital game design [32], and filmmaking [30]. The mainstream approaches can be broadly categorized into two groups: optimization-based methods and feed-forward methods. Optimization-based 3D generation methods leverage Score Distillation Sampling (SDS) [20] to iteratively refine 3D objects based on textual or visual inputs, effectively distilling the rich prior knowledge embedded in large pre-trained diffusion models [24]. In contrast, feed-forward 3D generation approaches [19, 28, 11, 12, 36, 29, 5] generate 3D content in a single inference pass, significantly improving efficiency and making real-time applications more feasible.
While significant progress has been made in 3D generation research, a critical yet under-explored challenge remains—the controllability of 3D generation. Controllable generation seeks to integrate precise control signals, such as edges, sketches, depth maps, surface normals, and text prompts, as conditioning inputs. By incorporating these structured constraints, controllable 3D generation enhances the flexibility and reliability of the synthesis process, enabling more precise manipulation of the generated content.
To advance research in controllable 3D generation, MVControl [15] extends ControlNet, originally designed for conditioned 2D image generation, to a multi-view diffusion model [25]. It first predicts four-view images, which are then processed by a multi-view Gaussian reconstruction model [27] to generate the final 3D representation. In contrast, ControLRM [31] aims to develop an end-to-end controllable 3D generation framework by replacing the original image encoder of a large reconstruction model (LRM) with a conditional encoder, enabling more precise conditioning. While these studies have demonstrated a degree of controllability, a significant challenge remains, i.e., achieving precise and fine-grained conditional control. As illustrated in Fig. 1, the extracted signals from the 3D models generated by MVControl and ControLRM exhibit noticeable discrepancies from the original input controls, underscoring the need for more refined and accurate controllable 3D generation.
To address this challenge, we propose CyC3D, a novel and general training framework that enhances the controllability of 3D generation. Our approach extends a feed-forward 3D generation network into a cyclic paradigm, enabling more precise and consistent adherence to input control signals, as illustrated in Fig. 2. Specifically, given an input control condition on a reference view, the generated 3D content is first rendered from randomly sampled new viewpoints. From these rendered images, the corresponding control conditions are extracted and fed back into the generation model as inputs, ensuring that the model learns to maintain control consistency across views. Subsequently, the model regenerates the 3D content, which is then rendered back to the initial reference viewpoint, where the extracted control conditions should match the original input condition. To further refine controllability, we introduce two complementary consistency constraints: View Cycle Consistency, which ensures that the two generated contents have similar semantics in embedded space since they are generated with the same prompt and highly coupled conditions, and Condition Cycle Consistency, which enforces strict alignment between the extracted control signals from the final rendered images and the original input condition. We compare our CyC3D against state-of-the-art approaches on GSO and ABO datasets [31] for zero-shot evaluation. Experimental results demonstrate that our framework significantly enhances controllability across all benchmarks while preserving competitive generation quality compared to existing methods.
We highlight the contributions of this work as follows:
  • New Insight and Framework: We reveal that existing efforts in controllable 3D generation still perform poorly on fine-grained controllability, and propose a novel framework (CyC3D) to handle the problem, which extends the original single-stage novel view synthesis pipeline to a double-stage one with cycle consistency (Fig. 2).
  • Cyclic Consistency Feedback: To refine controllability, we propose two complementary consistency constraints: condition cycle consistency and view cycle consistency. The former regularizes the consistency between input and reconstructed conditions, the latter enhances the semantic consistency of different stages.
  • Evaluation and Promising Results: We provide a comprehensive zero-shot evaluation of the controllability on GSO/ABO benchmarks, and demonstrate that CyC3D comprehensively outperforms existing methods.
2 Related Work
Text-/Image-to-3D Generation. 3D generation can be broadly categorized into optimization-based and feed-forward approaches: (1) Optimization-based methods offer a viable alternative to traditional techniques by circumventing the need for extensive text-to-3D datasets. DreamFusion [20] exemplifies this approach by utilizing SDS loss to optimize neural fields with diffusion priors, resulting in high-quality 3D assets. Despite their effectiveness, these methods often encounter the Janus problem, which arises from inconsistencies between 2D supervision and 3D geometry. (2) Feed-forward models have emerged as a prominent approach for efficient 3D generation, leveraging extensive 3D datasets. Techniques such as LRM [12] and TripoSR [29] utilize transformer architectures to predict NeRF from single-view images, demonstrating high generalization capabilities. Multi-view methods extend these approaches by synthesizing higher-quality 3D models from single-view images using multi-view diffusion models, as demonstrated by SyncDreamer [16] and MVDream [25].
Controllable 3D Generation. Achieving controllability in 3D generation remains a critical objective in the field. Techniques developed for controllable 2D generation have provided valuable insights and methods that are applicable to the more complex domain of controllable 3D generation. Inspired by ControlNet [35], MVControl [15] integrates a trainable control network with a multi-view diffusion model to facilitate controllable multi-view image generation. This method combines coarse Gaussian generation with SDS-based refinement but faces challenges related to model generalization and procedural complexity. ControLRM [31] addresses these limitations with its end-to-end feed-forward model, ensuring rapid inference and enhanced control precision, thus offering an efficient and effective solution for controllable 3D generation. Building on the advances of ControlNet++ [14] and ControLRM [35], we improve conditional controls to achieve more sophisticated and nuanced control over the generated 3D assets.
3 CyC3D
Figure 3: Overall framework of our proposed CyC3D. (a) summarize the procedure of our CyC3D. (b) shows the condition extractor utilized for different conditional controls (Sec. 3.2). (c) illustrates the condition cycle consistency regularization (Sec. 3.3). (d) introduces the proposed view cycle consistency regularization (Sec. 3.4).
As shown in Fig. 3, we present the overall framework of our proposed CyC3D. In Sec. 3.1, we first introduce the feed-forward backbone network for controllable 3D generation in our CyC3D. Then we introduce the condition extractor (Fig. 3(b)) for different conditional controls (e.g. edge, sketch, depth, normal) in Sec. 3.2. In Sec. 3.3, we introduce the design of condition cycle consistency regularization (Fig. 3(c)) to enhance the fined-grained controls. In Sec. 3.4, we further introduce the proposed view cycle consistency regularization (Fig. 3(d)) to stabilize the training and avoid overfitting on the input condition. Finally, we introduce the overall loss in Sec. 3.5.
3.1 Generation Backbone
For controllable 3D generation, our CyC3D can be combined with arbitrary feed-forward backbones (e.g. ControLRM [31], MVControl [15]). In default, we select ControLRM [31] as our backbone model. ControLRM contains a 2D condition extractor and a 3D triplane decoder. (1) The conditional generator in ControLRM is designed to take both text and 2D visual conditions as inputs. This generator can utilize either a transformer or diffusion backbone to create initial 2D latent representations, which are subsequently transformed into 3D models. (2) The 3D triplane decoder is inherited from a pre-trained large reconstruction model [12]. The triplane transformer decoder employs a unique representation of 3D data by decomposing it into three orthogonal planes (X, Y, and Z). Each plane is processed using transformer layers, which utilize self-attention mechanisms to capture spatial relationships and dependencies within the data.
3.2 Condition Extractor
To ensure the gradient back-propagation of cycle consistency in our CyC3D framework during training, the utilized condition extractor in Fig. 2 should be differentiable. The design of the condition extractor for each condition control is shown as follows:
Canny Condition. As shown in the figure, we extract edge maps using the Canny algorithm [4]. Following [14], a differentiable Canny extractor can implemented based on Kornia [23] to extract edge maps.
Sketch Condition. Following ControlNet [35], we utilize a pre-trained CNN network to extract sketch conditions from images. Since the CNN network is differentiable, the gradients can be passed backward through the network and back-propagated to the remaining modules.
Depth Condition. In analogy with the sketch condition, we also utilize a pre-trained foundation model (Depth Anything [33]) to extract the depth condition from rendered images. The whole process is differentiable in the cycle consistency pipeline.
Normal Condition. In practice, we find that the existing normal estimation models [1] utilized in ControlNet [35] tends to predict over-smoothed normal maps. It may result in a large gap between the exact 3D shape of the rendered multi-view images and the estimated normal maps. Consequently, we adopt an alternative solution to predict the depth map first and then convert the depth map to normal map via a differentiable module [13]. The great zero-shot generalization ability of large depth-estimation models like Depth-Anything [33] can effectively capture the 3D prior from the image and thus generate accurate normal map.
3.3 Condition Cycle Consistency Regularization
Denote the input edge condition as $C_{e}\in\mathbb{R}^{H\times W}$, the sketch condition as $C_{s}\in\mathbb{R}^{H\times W}$, the depth condition as $C_{d}\in\mathbb{R}^{H\times W}$, and the normal condition as $C_{n}\in\mathbb{R}^{H\times W\times 3}$. The input visual condition is noted as $C_{v}\in\{C_{e},C_{s},C_{d},C_{n}\}$. Given the text prompt $C_{t}$ and the input visual condition as $C_{v}$, the generation model outputs the triplane representation $P\in\mathbb{R}^{3\times C_{p}\times H_{p}\times W_{p}}$ corresponding to the input conditions.
$\small P_{i}=f_{\text{gen}}(C_{v},C_{t},T_{i}),$ (1)
where $T_{i}$ is the camera pose of the input condition, which is usually assumed to be an identity matrix after normalizing all the camera extrinsics to a frontal view. $f_{\text{gen}}$ is the backbone generation model for controllable 3D generation.
Given a 3D point $x\in\mathbb{R}^{3}$, the color and density filed $(c,\sigma)$ can be obtained via sampling the corresponding features on the triplane $P_{i}$: $(c,\sigma)=f_{\text{mlp}}(P_{i},x)$. By sampling the points along the viewpoint $T_{r}\in\mathbb{R}^{4\times 4}$, we can render the image on view $T_{r}$ via volume rendering [17]:
$\small\hat{I_{r}},\hat{D_{r}}=f_{\text{render}}(P_{i},T_{r}),$ (2)
where $T_{r}$ is the randomly sampled viewpoints as shown in Fig. 3. $\hat{I_{r}}$ is the rendered image on view $T_{r}$, and $\hat{D_{r}}$ is the rendered depth map. Then we can use the corresponding condition extractor $f_{\text{cond}}$ introduced in Sec. 3.2 and Fig. 3 (b) to extract the condition map on the sampled view $T_{r}$:
$\small\hat{C_{r}}=f_{\text{cond}}(\hat{I_{r}}),$ (3)
where $\hat{C_{r}}$ is the extracted condition map. Note that the condition extractor $f_{\text{cond}}$ is differentiable.
Figure 4: Visualization results of our CyC3D. The input conditions and the rendered images at different views are visualized. Our CyC3D can generate high quality results with consistent conditional controls.
In a second time, we can feed the obtained condition map $\hat{C_{r}}$ on novel view $T_{r}$ back to the generation model $f_{\text{gen}}$ again:
$\small P_{r}=f_{\text{gen}}(\hat{C_{r}},C_{t},T_{r}),$ (4)
where the predicted triplane $P_{r}$ can further be used to render on the input reference view $T_{i}$:
$\small\overline{I_{i}},\overline{D_{i}}=f_{\text{render}}(P_{r},T_{i}),$ (5)
where $\overline{I_{i}}$ is the rendered image on the input reference view, and $\overline{D_{i}}$ is the rendered depth map. Then we can extract the conditon map from the rendered image $\overline{I_{i}}$:
$\small\overline{C_{i}}=f_{\text{cond}}(\overline{I_{i}}),$ (6)
where $\overline{C_{i}}$ is the generated condition map back on the input reference view.
As shown in Fig. 2 and 3, we can calculate the difference between the reconstructed condition and the input condition to regularize the fine-grained controls. Depending on the type of conditional controls, we also propose two different kinds of conditional cycle-consistency feedbacks as follows:
2D Conditional Control Feedback: If we only consider 2D conditional controls extracted from the rendered images (e.g. Canny map extracted from RGB image), we can use a condition extractor to reconstruct the control signal. The generated condition map $\overline{C_{i}}$ via Eq. 6 should be consistent with the original condition map $C_{v}$ on the input reference view.
$\small L_{\text{cond}}=\sum_{T_{r}}\|\overline{C_{i}}-C_{v}\|_{2}^{2}.$ (7)
If the randomly sampled camera pose $T_{r}$ equals to the input viewpoint $T_{i}$, the two-step inference in cycle consistency will be degraded to a single one (as shown in Fig. 3 (c)), saving the computation cost:
$\small\hat{I_{i}},\hat{D_{i}}=f_{\text{render}}(f_{\text{gen}}(C_{v},C_{t},T_{i}),T_{i}),$ (8)
$\small\hat{C_{i}}=f_{\text{cond}}(\hat{I_{i}}),$ (9)
where $\hat{C_{i}}$ is the extracted condition map on the input reference view. In this way, we can modify Eq. 7 as:
$\small L_{\text{cond}}=\sum_{T_{r}\neq T_{i}}\|\overline{C_{i}}-C_{v}\|_{2}^{2}+\|\hat{C_{i}}-C_{v}\|_{2}^{2}.$ (10)
Figure 5: Qualitative comparison among our CyC3D and the other state-of-the-art controllable 3D generation methods.
3D Conditional Control Feedback: If we consider 3D conditional controls like depth or normal conditions, the 3D shape of the generated objects should be regularized as well for a fine-grained control.
(1) If the input condition is depth map ($C_{v}=C_{d}$), the explicit consistency on the rendered depth map and the input conditional depth map can be calculated as follows:
$\small L_{\text{cond-d}}=\sum_{T_{r}}\|f_{\text{norm}}(\overline{D_{i}})-C_{d}\|_{2}^{2},$ (11)
where $f_{\text{norm}}$ is the depth normalizing function. Since the rendered depth $\overline{D_{i}}$ and the input depth $C_{d}$ have a gap in the scale of depth, we utilize $f_{\text{norm}}$ to normalize them to the same scale and calculate a scale-agnostic loss following [22].
In analogy with Eq. 10, when the randomly sampled camera pose $T_{r}$ equals to the input viewpoint $T_{i}$, there is no need to run the inference twice during propagation. We can modify Eq. 11 as follows:
$\small L_{\text{cond-d}}=\sum_{T_{r}\neq T_{i}}\|f_{\text{norm}}(\overline{D_{i}})-C_{d}\|_{2}^{2}+\|f_{\text{norm}}(\hat{D_{i}})-C_{d}\|_{2}^{2},$ (12)
where $\hat{D_{i}}$ can be obtained from Eq. 8.
(2) If the input condition is normal map ($C_{v}=C_{n}$), the explicit depth-normal consistency can be further calculated as follows:
$\small L_{\text{cond-n}}=\sum_{T_{r}}\|f_{\text{d2n}}(\overline{D_{i}})-C_{n}\|_{2}^{2},$ (13)
where $f_{\text{d2n}}$ is a differentiable module [13] that can converts the depth map to a unified normal map.
In analogy with Eq. 10 and 12, when the randomly sampled camera pose $T_{r}$ equals to the input viewpoint $T_{i}$, there is no need to run the inference twice during propagation. We can thus modify Eq. 13 as follows:
$\small L_{\text{cond-n}}=\sum_{T_{r}\neq T_{i}}\|f_{\text{d2n}}(\overline{D_{i}})-C_{n}\|_{2}^{2}+\|f_{\text{d2n}}(\hat{D_{i}})-C_{n}\|_{2}^{2},$ (14)
where $\hat{D_{i}}$ can be obtained from Eq. 8.
Figure 6: Qualitative comparison of controllability among our CyC3D and the existing state-of-the-art methods in controllable 3D generation.
3.4 View Cycle Consistency Regularization
Since the condition cycle consistency regularization (Sec. 3.3) is a strong constraint that enforces the reconstructed and the input conditions to be as similar as possible, only using this loss might lead to overfitting towards the input viewpoint. If the generated 3D content from the initial view fails, the remaining processes might be misled to a wrong cross-view mapping. Only using the condition cycle consistency on the input view can not handle this issue, and the semantic multi-view consistency might be ignored during optimization. Consequently, we propose View Cycle Consistency Regularization to ensure convergence.
Rendering Regularization Loss: To regularize the possible failures in generation, the image reconstruction loss between the rendered images from generated 3D contents and the ground truth images on different views are used to supervise the generation model $f_{\text{gen}}$:
$\small L_{\text{render}}=\sum_{T_{l}}\|f_{\text{render}}(f_{\text{gen}}(C_{v},C_{i},T_{i}),T_{l})-I_{l}^{\text{gt}}\|_{2}^{2},$ (15)
where $T_{l}$ means the pre-defined ground truth viewpoints in the multi-view dataset, and $I_{l}^{\text{gt}}$ is the corresponding ground truth image. Note that this constraint can only be used in the first-time feedforward computation of the generation model. Because the randomness of the generation model might be depressed if enforcing the first-time and second-time generated 3D contents to be the same.
Semantic Consistency Loss: We utilize the text prompt $C_{t}$ to regularize the semantic consistency of different stages.
$\small L_{\text{clip}}=\sum_{T_{r}}(1-\cos(f_{\text{clip-t}}(C_{t}),f_{\text{clip-i}}(\hat{I}_{r})))+$ (16)
$(1-\cos(f_{\text{clip-t}}(C_{t}),f_{\text{clip-i}}(\overline{I}_{i}))).$
View Consistency Regularization: The overall view cycle consistency loss can be computed as follows:
$\small L_{\text{view}}=L_{\text{render}}+\alpha\cdot L_{\text{clip}},$ (17)
where $\alpha$ is the set to 5.0 following [31] to balance the scale of different terms as default.
Methods Edge Sketch Depth Normal
PSNR SSIM MSE PSNR SSIM MSE M-MSE Z-MSE R-MSE NB-MSE DN-CON
GSGEN [6] 13.01 0.780 0.0526 15.07 0.772 0.0322 0.1277 0.1268 0.0311 0.0230 0.0497
GaussianDreamer [34] 12.41 0.766 0.0597 14.58 0.792 0.0362 0.0740 0.0947 0.0391 0.0225 0.0366
DreamGaussian [26] 9.94 0.686 0.1048 14.42 0.765 0.0380 0.0980 0.0916 0.0406 0.0216 0.0335
VolumeDiffusion [28] 13.25 0.818 0.0497 16.50 0.836 0.0231 0.1202 0.0953 0.0531 0.0190 0.0172
3DTopia [11] 9.63 0.704 0.1134 15.54 0.801 0.0287 0.1193 0.1064 0.0420 0.0268 0.0372
MVControl [15] 10.44 0.723 0.0952 14.51 0.746 0.0370 0.0555 0.0645 0.0160 0.0209 0.0315
ControLRM-T [31] 12.43 0.836 0.0607 17.17 0.836 0.0202 0.0567 0.0713 0.0599 0.0120 0.0359
ControLRM-D [31] 12.47 0.837 0.0600 17.23 0.837 0.0197 0.0383 0.0546 0.0650 0.0119 0.0357
CyC3D-T (Ours) 15.17 0.885 0.0342 18.31 0.889 0.0159 0.0051 0.0037 0.0065 0.0027 0.0062
CyC3D-D (Ours) 15.09 0.885 0.0347 18.38 0.890 0.0156 0.0051 0.0041 0.0060 0.0028 0.0060
Table 1: Quantitative comparison of controllability with state-of-the-art controllable 3D generation methods on GSO benchmark. $\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better. The best results are highlighted in bold. We provide the results of four different conditions: edge, sketch, depth, and normal. Our CyC3D-T/-D achieves significant controllability improvements in the table.
Methods Edge Sketch Depth Normal
PSNR SSIM MSE PSNR SSIM MSE M-MSE Z-MSE R-MSE NB-MSE DN-CON
GSGEN [6] 12.75 0.737 0.0614 14.25 0.744 0.0384 0.1004 0.0967 0.0463 0.0171 0.0479
GaussianDreamer [34] 11.92 0.736 0.0699 13.75 0.734 0.0439 0.0809 0.1061 0.0521 0.0188 0.0463
DreamGaussian [26] 9.38 0.683 0.1178 12.90 0.716 0.0523 0.1267 0.0762 0.0433 0.0188 0.0349
VolumeDiffusion [28] 12.58 0.734 0.0619 15.03 0.780 0.0321 0.1007 0.1198 0.0671 0.0188 0.0264
3DTopia [11] 8.48 0.633 0.1469 14.41 0.741 0.0372 0.1916 0.1844 0.0559 0.0346 0.0418
MVControl [15] 9.52 0.656 0.1161 13.34 0.686 0.0479 0.0454 0.0467 0.0181 0.0133 0.0350
ControLRM-T [31] 10.49 0.552 0.0909 15.77 0.777 0.0272 0.0514 0.0497 0.0700 0.0136 0.0572
ControLRM-D [31] 11.51 0.592 0.0725 15.80 0.778 0.0269 0.0493 0.0545 0.0661 0.0158 0.0548
CyC3D-T (Ours) 12.98 0.756 0.0540 16.55 0.847 0.0225 0.0080 0.0078 0.0102 0.0029 0.0102
CyC3D-D (Ours) 12.99 0.755 0.0539 16.45 0.845 0.0231 0.0075 0.0087 0.0091 0.0030 0.0109
Table 2: Quantitative comparison of controllability with state-of-the-art controllable 3D generation methods on ABO benchmark. $\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better. The best results are highlighted in bold. We provide the results of four different conditions: edge, sketch, depth, and normal. Our CyC3D-T/-D achieves significant controllability improvements in the table.
3.5 Overall Loss
Finally, the overall loss is the combination of the aforementioned losses:
$L_{\text{total}}=\left\{\begin{aligned} &L_{\text{view}}+\lambda\cdot L_{\text{cond}}+\beta\cdot L_{\text{cond-d}},\;\;\text{if}\;\;C_{v}=C_{d}\\ &L_{\text{view}}+\lambda\cdot L_{\text{cond}}+\beta\cdot L_{\text{cond-n}},\;\;\text{if}\;\;C_{v}=C_{n}\\ &L_{\text{view}}+\lambda\cdot L_{\text{cond}},\quad\text{otherwise}\end{aligned}\right.$ (18)
where $\lambda$ is set to 1.0 in default, and $\beta$ is set to 0.1 for regularization.
4 Experiment
4.1 Implementation Details
Datasets. Our training dataset is the training split of G-Objaverse (GOBJ) dataset [21], a sub-set of Objaverse [8]. Following the selection principle of [15], we utilize the test samples collected by [31] gathered from 2 different datasets for zero-shot evaluation: 80 samples from Google Scanned Objects dataset (GSO) [9], and 80 samples from Amazon Berkeley Objects dataset (ABO) [7].
[…]
FID CLIP-I CLIP-T FID CLIP-I CLIP-T
GSGEN [6] 344.61 0.740 0.289 366.47 0.669 0.259
GaussianDreamer [34] 278.70 0.810 0.300 225.38 0.787 0.277
DreamGaussian [26] 359.65 0.760 0.279 392.95 0.723 0.247
VolumeDiffusion [28] 299.61 0.715 0.259 350.46 0.679 0.242
3DTopia [11] 331.39 0.727 0.283 231.55 0.751 0.272
MVControl [15] 278.08 0.816 0.288 217.97 0.802 0.291
ControLRM-T [31] 260.75 0.846 0.289 202.14 0.827 0.282
ControLRM-D [31] 171.13 0.838 0.302 181.84 0.836 0.292
CyC3D-T (Ours) 255.30 0.840 0.295 196.55 0.836 0.291
CyC3D-D (Ours) 168.71 0.846 0.303 183.84 0.842 0.293
Table 3: Generation quality (FID) and CLIP score (CLIP-I, CLIP-T) comparison with state-of-the-art controllable 3D generation methods GSO, and ABO benchmarks. $\uparrow$ means higher result is better, while $\downarrow$ means lower is better. The best results are emphasized in bold. We provide the average results of 4 different conditions (edge, sketch, depth, and normal) in the table.
Training Details. In analogy with [31], we initialize our network with the weights from pretrained OpenLRM-base [12]. For CyC3D-T, the conditional backbone is a transformer-based network; For CyC3D-D, the conditional backbone is a diffusion-based network. The training might cost 2 days for CyC3D-T and 3 days for CyC3D-D on 32 Nvidia V100-32G GPUs. More details are provided in the appendix.
Evaluation and Metrics. For the task of controllable 3D generation task, we consider 4 different conditions: edge, sketch, depth, and normal. For each condition, we evaluate the controllability by measuring the similarity between the input conditions and the extracted conditions from generated 3D contents following [31]. For edge/sketch condition, we utilize PSNR, SSIM, and MSE as evaluation metrics; For depth/normal condition, we utilize MSE to measure the similarity. In analogy with [14], the large foundation models in monocular depth estimation are used to capture the depth priors for model-based evaluation. If Midas [22] is used, the metric is noted as M-MSE. If ZoeDepth [3] is used, the metric is noted as Z-MSE. Following [31], the depth consistency in 3D space compares the difference between rendered depth map and the input condition depth map, noted as R-MSE. In analogy with [14], the metric of NB-MSE uses the pre-trained model (Normal-BAE [1]) in surface normal estimation to assess the normal consistency. Furthermore, we also evaluate the depth-normal consistency between the rendered depth map and the input normal maps following [31], noted as DN-CON. Please refer to more detailed introduction of these evaluation metrics in the appendix.
4.2 Experimental Results
Methods Control Quality
Edge Sketch Depth Normal FID CLIP-I CLIP-T
ControlNet+OpenLRM-S [12] 0.1069 0.0459 0.1333 0.0255 268.7 0.807 0.292
ControlNet+OpenLRM-B [12] 0.0897 0.0452 0.1312 0.0206 258.04 0.822 0.293
ControlNet+OpenLRM-L [12] 0.0792 0.0447 0.1383 0.0226 256.43 0.824 0.294
ControlNet+TGS [36] 0.0523 0.0417 0.1403 0.0208 258.86 0.841 0.295
ControlNet+TripoSR [29] 0.0978 0.0421 0.1188 0.0241 261.76 0.835 0.289
MVControlNet+LGM [15] 0.0952 0.0370 0.0160 0.0315 278.08 0.816 0.288
CyC3D-T (Ours) 0.0342 0.0159 0.0065 0.0062 255.30 0.840 0.295
CyC3D-D (Ours) 0.0347 0.0156 0.0060 0.0060 168.71 0.846 0.303
Table 4: Comparison with ControlNet-based baselines (ControlNet + Image-to-3D methods) on GSO benchmark.
Qualitative Results. We provide qualitative results of our CyC3D in Fig. 4. The backbone is inherited from ControLRM [31], and the diffusion model is replaced with a pre-trained ControlNet. As the figure shows, our method not only performs well in generating hight quality 3D contents but also in preserving details in conditional controls. Furthermore, we also compare the qualitative results of our CyC3D with other state-of-the-art methods including ControLRM [31] and MVControl [15] in Fig. 5.
Controllability Evaluation. The comparisons of controllability with state-of-the-art controllable 3D generation methods on GSO/ABO benchmarks are presented in Tab. 1/2. Following [31], we report the controllability metrics discussed in Sec. 4.1 on all conditional controls. From the tables, we can find that our CyC3D can achieve state-of-the-art performance in the quantitative comparison with other baselines. For example, the R-MSE of our CyC3D-D is 0.0060, which is much lower than the baselines of ControLRM-D (0.0650 R-MSE). Furthermore, we also provide qualitative comparison of controllability in Fig. 6.
Generation Evaluation. To investigate whether improving controllability compromises generation quality, we provide the quantitative results of Fréchet Inception Distance (FID) [10] and CLIP-score [18] following [15] in Tab. 3. We report CLIP-I metric to quantify image similarity between reference and rendered images, and CLIP-T metric to evaluate text-image alignment using textual annotations. In rigorous zero-shot generalization experiments of cross-domain transitions towards GSO and ABO, our proposed CyC3D demonstrates statistically significant improvements over existing methods. Furthermore, from the qualitative results in Fig. 5, out CyC3D can generate better results with higher quality and consistency than existing methods.
Comparison with ControlNet-based Baselines. For a fair comparison, we also provide experimental results compared with ControlNet combined with image-to-3D reconstruction models. The images produced by ControlNet are processed through image-to-3D reconstruction models to synthesize 3D assets, with foreground segmentation implemented using Rembg for background removal. The quantitative evaluation results on the GSO benchmark are presented in Tab. 4, incorporating state-of-the-art image-to-3D systems including OpenLRM [12], TGS [36], and TripoSR [29] as comparative baselines.
Loss Edge Sketch
PSNR SSIM MSE PSNR SSIM MSE
w/o $L_{\text{cond}}$ 12.87 0.841 0.0570 17.52 0.888 0.0187
w $L_{\text{cond}}$ 15.09 0.885 0.0347 18.38 0.890 0.0156
Table 5: Ablation experiments of 2D condition cycle consistency.
Loss Depth Loss Normal
M-MSE Z-MSE R-MSE NB-MSE DN-CON
w/o $L_{\text{cond}}$ 0.0383 0.0546 0.0650 w/o $L_{\text{cond}}$ 0.0034 0.0205
w $L_{\text{cond}}$ 0.0192 0.0241 0.0321 w $L_{\text{cond}}$ 0.0023 0.0187
w $L_{\text{cond}}+L_{\text{cond-d}}$ 0.0051 0.0041 0.0060 w $L_{\text{cond}}+L_{\text{cond-n}}$ 0.0011 0.0056
Table 6: Ablation experiments of 3D condition cycle consistency.
Figure 7: Ablation experiments of view cycle consistency.
4.3 Ablation Study
Effect of Condition Cycle Consistency. To evaluate the effectiveness of 2D conditional control feedback (e.g. Eq. 10 for edge/sketch), we provide ablation experiments in Tab. 5. Furthermore, we also provide ablation results of 3D conditional control feedback (Eq. 12 for depth, Eq. 14 for normal) in Tab. 6. As the tables show, each term of condition cycle consistency can effectively imporve the controllability upon different conditional controls.
Effect of View Cycle Consistency. To evaluate the effectiveness of view cycle consistency feedback (e.g. Eq. 17), we provide qualitative ablation results in Fig. 7. As the figure shows, without view cycle consistency regularization, the generated 3D contents might be overfitting to the input viewpoint. Though the rendered results in the input view is reasonable, the lack of regularization on the remaining views might lead to a low-quality 3D shape.
5 Conclusion
In this paper, we propose CyC3D, a novel framework for improving the controllability during 3D generation guided by different conditional controls (edge/sketch/depth/normal). The core idea is the cycle-view consistency which ensures accurate alignment between input conditions and generated models by rendering the generated 3D content from various viewpoints and re-extracting the conditions from these rendered views to feed back into the generation process. The quantitative and qualitative evaluation results on GSO and ABO benchmarks show the state-of-the-art controllability of CyC3D.